1.3 Minimize z = a + b + c
Subject to:
1. a - b - c ≤ 0
2. a + b + c ≥ 4
3. a + b - c = 2
4. a, b ≥ 0
1
Expert's answer
2020-04-09T14:44:49-0400

Clearly; min(z)=a+b+c=4

as it is given that "a+b+c \geq4"

One of the feasible solution is (a,b,c)=(1,2,1)

Checking this with the given constraints, we get;

1. "a - b - c ≤ 0 \implies 1-2-1=-2≤0"

2. "a + b + c ≥ 4 \implies 1+2+1=4≥ 4"

3. "a + b - c =2 \implies 1+2-1= 2=2"

4. "a, b ≥ 0 \implies 1 ≥ 0, 2 ≥ 0"

As all the constraints are satisfied, the proposed solution is valid.